OpenStudy (anonymous):

The effect of your toss is that the package is moving upward at 10.5 m/s as it leaves your hand. If the 2.5 kg package started at rest and the tossing action lasted 0.39 seconds, what was the net force acting on the package while you were tossing it and what was the strength of the force exerted on the ball by your hand?

OpenStudy (anonymous):

@matt101

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (alexrobin13):

First find acceleration. delta a = delta v / delta t (delta means the difference essentially) a = (10.5-0)/0.39 a=25.6m/s^2 F=ma F=(2.5)(25.6) F=64N

OpenStudy (anonymous):

Okay so is that how you find the net force or the force from the hand?

OpenStudy (badhi):

so the resultant force which was found to be 64N is a summation of the weight and the force by the hand  \(F_h-mg = 64N\)

OpenStudy (anonymous):

Okay so this is how you solve for Fnet or the force from the hand?

OpenStudy (anonymous):

Because Fh = 39N since you would take (-9.81*2.5)+64

OpenStudy (badhi):

the final force from the weight and the hand combined is equal to 64N because eventually that resultant force is what generates the acceleration. this 64N is generated by (as i previously told) from the force and the weight (\(F_h-mg\) to be precise) so \(F_h = mg+64N\) (note that \(F_h\) should be more than 64N )

OpenStudy (anonymous):

Okay so the Fh would be 88.5 or 89. So would you take the 89N and subtract it from the 64N to get the net force?

OpenStudy (anonymous):

Also sorry about the late reply. I was working on a paper and didn't notice you replied.

OpenStudy (badhi):

net force is the vector summation of the \(F_h\) and \(mg\) which is equal to 64N that causes an acceleration of 25.6ms^(-2)

OpenStudy (anonymous):

Okay so I should add the 89N with the 64N to get net force? Is the 89N even right?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Okay for the first bit do you take the 10.5 and add it to the 26.92?

OpenStudy (unklerhaukus):

The first step to to find the acceleration of the package, we know the change in velocity, and the change in time, the acceleration is the ratio of these: \[a= \frac{\Delta v}{\Delta t}\] What do you get for this?

OpenStudy (anonymous):

Well the Vf would be 10.5m/s and it started from rest so the Vi would be 0. The t is .39s right?

OpenStudy (anonymous):

So (10.5-0)/.39

OpenStudy (unklerhaukus):

yeah, . .

OpenStudy (anonymous):

And get the 26.92 m.s^2

OpenStudy (unklerhaukus):

that's right, but the units are m/s^2

OpenStudy (anonymous):

Sorry. typo

OpenStudy (unklerhaukus):

Now we can calculate the net force on the package, with the mass. \[F_\text{net}= ma =\] What do you get for this?

OpenStudy (anonymous):

Fnet = 2.5kg * 26.92m/s^2

OpenStudy (anonymous):

=67.31N

OpenStudy (unklerhaukus):

good, The net force includes the force of gravity, (which is acting in the opposite direction to the toss) so the force of the toss must be greater than the net force, by the force of gravity \[F_\text{toss}=F_\text{net}+F_\text{gravity}\] where gravitational acceleration \(g= 9.81\,[\text{m/s}^2]\), (assuming you are on earth)

OpenStudy (anonymous):

57.50N

OpenStudy (unklerhaukus):

the force of the toss should be greater than the net force

OpenStudy (anonymous):

So you don't take gravity as a negative?

OpenStudy (unklerhaukus):

\[\textbf F_\text{toss}+\textbf F_\text{gravity}=\textbf F_\text{net}\\ F_\text{toss}\hat{\textbf y}+ F_\text{gravity}(-\hat{\textbf y})= F_\text{net}\hat{\textbf y}\\ F_\text{toss}- F_\text{gravity}= F_\text{net}\\ F_\text{toss}= F_\text{net}+ F_\text{gravity}\\ \]

OpenStudy (anonymous):

What is the -y?

OpenStudy (unklerhaukus):

the gravitational force vector points down, the toss went up

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So should I get 77.12N then?

OpenStudy (unklerhaukus):

what is the magnitude of gravitational force on a 2.5kg object?

OpenStudy (anonymous):

No idea

OpenStudy (unklerhaukus):

use F = ma, where a = g = 9.81 [m/s^2], m = 2.5 [kg]

OpenStudy (anonymous):

Do 24.53

OpenStudy (unklerhaukus):

yeah

OpenStudy (anonymous):

*So

OpenStudy (unklerhaukus):

So \[F_\text{toss}= F_\text{net}+ F_\text{gravity}\\ \qquad= 67.31\,[\text N]+24.52\,[\text N]\\ \qquad=\]

OpenStudy (anonymous):

91.84N

OpenStudy (unklerhaukus):

Good.

OpenStudy (unklerhaukus):

Is the ball the same thing as the package?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Well I guess it would be

OpenStudy (anonymous):

Okay so the Fnet = 67.31N and the Fhand = 91.84N?

OpenStudy (unklerhaukus):

i think so,

OpenStudy (anonymous):

So it would be two sig fig since the mass had the least and it was two right?

OpenStudy (unklerhaukus):

good point.

OpenStudy (anonymous):

Awesome. I got it right. Thank you!!!

OpenStudy (anonymous):

I've been curious about this. Is your username from the show?

OpenStudy (unklerhaukus):

it isn't

OpenStudy (anonymous):

Alright. I've just always wondered that.

OpenStudy (anonymous):

Do you have time for another problem by chance? It's fine if you don't.